# Predictively Oriented Posteriors

### v1

@@ 5 Illustrations
We end our development on PrO posteriors by illustrating some of their key behaviours. First, we explore the practical ramifications of convex recovery and the choice of different scoring rule approximations in Section 5.1 in the context of classification. Next, we show that similar conclusions hold for the continuous case by constructing in Section 5.2, and provide a more thorough investigation into how PrO posteriors achieve better predictive performance than Gibbs posteriors. To conclude, we highlight the potential future use of PrO posteriors for diagnosing model misspecification in a real world data set in Section 5.3.
@@ 5.1 Binary classification
For a first assessment of the predictive benefits conferred by PrO posteriors, we consider a simple synthetic classification example in which the model allows for convex recovery as set out in Definition 2. Labels $y_{i}$ are assigned deterministically as a function of the two-dimensional covariate vector $(x_{i}^{(0)},x_{i}^{(1)})\in[-2,2]^{2}$. Throughout, $(x_{i}^{(0)},x_{i}^{(1)})$ are sampled uniformly. When the covariates fall into the top-left quadrant of $2,2]^{2}$ (so that $x_{i}^{(0)}<0,x_{i}^{(1)}>0$), $y_{i}=0$. In contrast, the bottom-left quadrant (for which $x_{i}^{(0)}>0,x_{i}^{(i)}<0$) yields $y_{i}=1$, while labels in the remaining two quadrants are drawn from a Bernoulli distribution with parameter $0.5$.
We fit these data with a logistic classifier, and show the results in Figure 5. The plot compares the predictive performance of the Bayes posterior with that of a PrO posterior based on an approximation to the predictive log score. [^1] Here, we slightly abuse terminology: the Bayes posterior in this section is in fact tempered by $\lambda_{n}=n^{1/2}$. While this scaling term is required for our theory to apply, the experimental evidence in this section and its interpretation would be the same if the Bayes posterior were not tempered and $\lambda_{n}=1$. Since this approximation can be obtained either by following the DI-based approach of Masegosa (2020) that adheres to part $\operatorname{(DI)}$ of Assumption 2 or the MS-based approach which satisfies part $\operatorname{(MS)}$ and was originally put forward by Morningstar et al. (2022), we plot both of them. From the pictorial evidence, it is clear that relative to either approximation scheme, the Gibbs posterior produces a subpar decision boundary, and leads to overconfident and incorrect predictions—particularly in the top-right and bottom-left quadrants. In contrast, the PrO posteriors identify the correct decision boundaries, and recover a predictive distribution $P_{Q_{n}}$ for which $\mathrm{KL}(P_{0}\|P_{Q_{n}})$ is close to zero. This mirrors the predictions of Theorems 1 and 2, and is a consequence of the fact that the PrO posterior predictive allows for convex recovery of the exact decision boundary. Notably, there is little practical difference between the PrO posteriors in this setting: the predictive performance is similar whether we use the DI-based approximation or an MS-based approximation for $k=2$. This reinforces one of the key points of Theorem 2: even if the predictive score has to be approximated with an approximation error that cannot be recovered, the resulting PrO posterior will still predictively dominate the Bayes posterior.
Figure 5: Binary classification example. The first row shows the posterior predictive distributions of the Bayes posterior and the PrO posterior using two different approximations of the predictive logarithmic scoring rule; and is overlaid with the raw data. The second row describes the modes of the posterior predictive distributions, which gives rise to their decision boundaries.
5.2 Linear Regression
The previous section compared Gibbs and PrO posteriors for discrete-valued data. The following example extends this analysis, and compares performance in the context of a continuously-valued response when the model is non-trivially misspecified and allows for convex recovery. To do so, we generate outcomes from a mixture of linear regressions with the two-dimensional covariate $x_{i}\in\mathbb{R}^{2}$ distributed as standard normal, and with Gaussian errors $\varepsilon_{i}\sim\mathsf{N}(0,\sigma^{2})$. In particular, we generate them as
$y_{i}=\begin{cases}x_{i}^{\top}(0,2)+\varepsilon_{i}&\text{with probability }0.5\\ x_{i}^{\top}(0,-2)+\varepsilon_{i}&\text{otherwise}.\end{cases}$
We then fit these data using an MMD-based Gibbs posterior and a PrO posterior using the exact predictive score. To this end, we follow the formulation of MMD in the context of regression formulated by Alquier and Gerber (2024). Additional details on this and data generation protocols are deferred to Appendices D.1 and D.6.
We summarise our findings in the three panels of Figure 6. The left panel compares the posterior densities on the parameter space. It shows that the Gibbs posterior concentrates onto some convex combination of the underlying regression coefficients given by $(\theta_{1},\theta_{2})=(0,-c\cdot 2+(1-c)\cdot 2)$ for some $c\in(0,1)$. In contrast, the PrO posterior seems to concentrate onto the actual coefficients $(\theta_{1},\theta_{2})\in\{(0,-2),(0,2)\}$ which generated the data. This difference translates directly into predictive performance: the middle panel illustrates how the overconfidence of the Gibbs posterior results in a predictive distribution that is under-dispersed. The PrO posterior does not share this issue, and provides excellent predictive fit. This difference in predictive performance is quantified in the right panel, which demonstrates that the Gibbs posterior predictive performs roughly as well as the PrO posterior for data drawn from the one mode it focuses on, and substantively worse for all data from the mode which it neglects. Further details are provided in Appendix D.6.
Figure 6: Synthetic linear regression data. We show, from left to right, the joint posterior distribution of the PrO posterior and of the Gibbs posterior overlaid with the true regression coefficients (in black crosses); draws from their respective posterior predictive distributions are overlaid with the empirical distribution of the test data (in grey); and the difference between the PrO and Gibbs log point-wise predictive density for each test point. If the difference is positive, the PrO posterior predictive at that point is more accurate.
5.3 Putting Data
We end the section by considering a real world data set which can help us to compare the behaviour of Bayes and PrO posteriors in the case of non-trivial misspecification. Gelman and Nolan (2002) present data on professional golf putting success as a function of the distance from which the shot is taken. We show the proportion of successful putts in Figure 7 and fit a logistic regression to these data. The top row of this plot shows that the Bayes posterior predictive is overconfident, which is due to a highly concentrated parameter posterior. In contrast, the PrO posterior predictive is much better calibrated: it adapts to the present level of model misspecification, and does not concentrate around a single point, but instead onto two modes; each representing a distinct class of putting ability. In related but distinct contexts, this notable difference in uncertainties has led previous authors to speculate that one may be able to use PrO posterior-like methods as a diagnostic tool for model misspecification (see Lai and Yao, 2024; Chazal et al., 2025). While previous work argued for this on the basis of intuition and heuristics, our findings in Section 3 provide a way forward that would allow us a rigorous way for using PrO posteriors in the detection of model misspecification.
Figure 7: Golf putting data. Points represent the proportion of successful putts made by professional golfers as a function of the distance from the hole. The first column shows the predictive mean and $80\%$ credible intervals of the PrO posterior and of the Gibbs posterior. The second and third columns show the marginals of the corresponding parameter posteriors.
@@ 6 Discussion
In this paper, we have advocated for a new statistical paradigm that combines the best aspects of both parameter inference and density estimation. This led to PrO posteriors, which can operate at a similar predictive accuracy as density estimation techniques while simultaneously delivering practically meaningful uncertainty quantification for interpretable model parameters. We argued that PrO posteriors bridge the two main statistical traditions identified by Breiman (2001), and asserted that they lead to a version of epistemological uncertainty that is more expansive, adapts to model misspecification, and can avoid posterior collapse whenever doing so is benefits prediction. Further, we proved that PrO posteriors predictively dominate both Bayes and Gibbs posteriors. After introducing sampling algorithms for PrO posteriors, we illustrated our findings on several numerical examples.
PrO posteriors are a rigorous exploration of themes that have been present in several research communities, and build directly on the pioneering work of Masegosa (2020), Morningstar et al. (2022), and Shen et al. (2025). In a more applied setting, related ideas have also previously been studied for variational methods (Futami et al., 2021; Futami et al., 2022; Wei et al., 2021; Lai and Yao, 2024), in the context of Gaussian processes (Jankowiak et al., 2020a, b), and for mixture models (Wang and Liu, 2019). Within statistical learning, an important area of research with significant parallels to PrO posteriors are PAC-Bayesian bounds which are concerned with introducing diversity into an ensemble. The most important example is weighted majority voting, whereby a set of classifiers are assigned weights, and the final prediction is the label supported by the largest total weight across the ensemble (Lacasse et al., 2006; Laviolette et al., 2011; Germain et al., 2015; Laviolette et al., 2017; Lorenzen et al., 2019; Masegosa et al., 2020; Viallard et al., 2021; Wu et al., 2021; Wu and Seldin, 2022; Vignault et al., 2023; Theisen et al., 2024). Beyond majority voting, related PAC-Bayesian theories were also discussed in Audibert (2004), Ortega et al. (2022), Hauptvogel and Igel (2024) and Pitas et al. (2025).
Throughout, robustness to model misspecification was an important motivation in our development of PrO posteriors. Robustness is also one of the key drivers behind the recent emergence of generalised Bayesian methods, and particularly Gibbs posteriors: for example, Altamirano et al. (2023a, b); Matsubara et al. (2022, 2023); Laplante et al. (2025); Duran-Martin et al. (2024) all propose Gibbs posteriors which produce posteriors that are Bayes-like, but robust to outliers. However, these methods guard against misspecification only if our main interest remains parameter inference and our misspecification can be described through the famous Huber contamination model (Huber, 1964). If the misspecification falls far outside of this class of model perturbations, Gibbs posteriors constructed with this setting in mind generally perform poorly. Though the underlying meta-model is more intricate than the Huber contamination model, a similar conclusion holds for the coarsened Bayes posteriors of Miller and Dunson (2019). In contrast, PrO posteriors do not explicitly rely on a meta-model: while the results in Theorems 1 and 2 demonstrate strong performance guarantees whenever the data-generating process is a mixture over elements in $\mathcal{M}_{\Theta}$, they also show that PrO posteriors continue to dominate the predictive outputs of Gibbs and Bayes posteriors well beyond this setting.
We presented our findings as a post-Bayesian idea: grounded in the type of epistemological uncertainty produced by Bayesian methods, but not reliant on Bayes’ rule. It is interesting to note that one could also interpret our methodology as a nonparametric frequentist estimation technique: for the case of the logarithmic scoring rule, one could even interpret the PrO posterior objective as an KL-regularised version of a nonparametric maximum likelihood estimation (e.g. Laird, 1978; Lindsay, 1995). Extending this logic, PrO posteriors for general scoring rules could be thought of as frequentist KL-regularised nonparametric minimum distance estimators for the model class $\{P_{Q}:Q\in\mathcal{P}(\Theta)\}$ parameterised by a measure $Q\in\mathcal{P}(\Theta)$. However, as our parameter of interest is actually $\theta\in\Theta$, and since PrO posteriors directly incorporate our prior beliefs, $\Pi$, we argue that they should foremost be viewed as belief distributions over the parameters of interest, and are therefore more closely related to Bayesian and generalised Bayesian methods.
Despite a thorough exploration of PrO posteriors in the current paper, much remains to be done. Perhaps the most pressing issue is computation: to improve the sampling scheme we proposed, more insight is needed on hyper-parameter tuning and convergence diagnostics. We are optimistic that the recent paper of Chazal et al. (2025) may remedy this: on top of providing a new suite of diagnostics and convergence guarantees, it also proposes new algorithms applicable to PrO posteriors. Beyond computation, there are several promising methodological avenues: for instance, we should be able to derive statistical tests that use PrO posteriors to diagnose model misspecification. Unlike previous tests, they would not operate on posterior summaries, but directly on the difference in uncertainty between PrO posteriors and Gibbs posteriors. Further, the current theory is limited to observations which are independently drawn from a common data-generating mechanism. As there is considerable interest in forecasting and prediction for time series, graphical models, and spatial settings, it would be of great practical interest to extend our findings to dependent data. While this would be a technically challenging undertaking, our experimental evidence suggests that PrO posteriors should be expected to perform well in these settings. Other questions that are currently unaddressed include the choice of scoring rule: in finite samples, are there settings where one scoring rule should be preferred over another? What are the differences between PrO posteriors and their posterior predictives as we vary the underlying scoring rules? Likewise, we have not discussed prior choice in the current paper: though it seems intuitive to think of the prior as being conceptually similar to the generalised Bayesian case, we have not attempted to formalise this analogy. Doing so would have the potential of shedding more light on the role of prior regularisation in PrO posteriors, and could raise several key questions. For example, can we quantify how exactly the prior impacts the PrO posterior? Similarly, is it possible to adapt the arguments given by Catoni (2007) to PrO posteriors, and to derive priors which are in some sense predictively optimal?
In conclusion then, PrO posteriors open up many new possibilities, and allow us to reason about parameter uncertainty and prediction-centric statistical methods in a new light. They mark a clear departure from existing paradigms focused on either parameter inference or density estimation, and introduce a novel form of epistemological uncertainty unaccounted for by classical statistical methods. While the introduction of PrO posteriors results in numerous challenges that must be resolved before they are ready for wide-spread adoption, we believe that the scale of these challenges is matched by their transformative potential for modern statistical science.

### v2

@@ 5 Illustrations
We supplement the discussion of the previous section by illustrating the promise of PrO posteriors on three statistical problems. In Section 5.1, we highlight the implications of convex recovery in a binary classification problem. Moving past independently sampled data, the remaining examples illustrate that PrO posteriors outperform Gibbs posteriors even when the theory’s assumptions are violated. In this vein, Section 5.2 focuses on predicting temporally dependent water flows of Stansted Brook in England, while Section 5.3 investigates the impact of using different scoring rules in a graphical model for spatially dependent house prices in Boston.
@@ 5.1 Binary classification
To strengthen intuitions about the predictive benefits conferred by PrO posteriors, we construct a synthetic data set in which a generalised linear model with logistic link function allows for convex recovery as set out in Definition 2. Labels $y_{i}$ are assigned deterministically as a function of the covariate vector $(x_{i}^{(0)},x_{i}^{(1)})$ which is sampled uniformly on $2,2]^{2}$. When the covariate vector falls into the top-left quadrant of $2,2]^{2}$ (so that $x_{i}^{(0)}<0,x_{i}^{(1)}>0$), $y_{i}=0$. In contrast, the bottom-left quadrant (for which $x_{i}^{(0)}>0,x_{i}^{(i)}<0$) yields $y_{i}=1$, while labels in the remaining two quadrants are sampled as $y_{i}\sim\operatorname{Ber}(0.5)$.
We generate $n=1,000$ observations from this model and compute both the corresponding Bayes posterior and the PrO posterior based on the logarithmic scoring rule. We compare their posterior predictive distributions in Figure 8, and conclude that Bayesian inference produces a subpar decision boundary which leads to an overconfident misinterpretation about the data-generating process in the top-right and bottom-left quadrants. In contrast, the PrO posterior identifies the correct decision boundary, and recovers a predictive distribution $P_{Q_{n}}$ for which $\mathrm{KL}(P_{0}\|P_{Q_{n}})$ is close to zero. This mirrors the predictions of Theorems 1, and is a consequence of the fact that the PrO posterior predictive allows for convex recovery of the exact decision boundary. Figure 9 further illustrates this by probing the effect of $n$, and confirms that the theory derived is practically relevant even for relatively small data set sizes.
Figure 8: Top row: Bayes and PrO posterior predictive distributions overlaid with the raw data. Bottom row: Decision boundaries induced by individual draws from the posterior predictives. While the Bayes posterior concentrates onto a single decision boundary, the PrO posterior is multimodal.
Figure 9: Constructing Bayes and PrO posteriors using $n$ observations, we plot their negative log likelihood (NLL) on test data, with standard error intervals obtained from 2020 replications.
5.2 River flow data
Throughout, our theoretical developments have used the assumption of independently sampled data. In practice however, prediction is often most interesting for sequentially or spatially dependent data. While the theoretical development provided in the current paper cannot account for this, the following example shows that PrO posteriors continue to function as expected in temporally dependent data (see also Shen et al., 2025).
Here, we focus on predicting the gauged daily flow of Stansted Brook north of London using measurements of the UK’s Environment Agency. We model the flow using an auto-regressive model of order seven with an intercept term, and compare the inferences obtained from a Gibbs posterior and a PrO posterior using the squared MMD. With this set-up, we fit the model using data from March to November 2002, and use it to obtain one-step-ahead predictions over the course of the transition from the exceptionally wet winter of 2002/2003 into the country’s driest February–-October period since 1921 (Marsh, 2004). This abrupt shift from heavy rain in January to the dry and warm conditions from February onwards produced severe though short-lived hydrological stress, and offers an interesting regime for testing the adaptivity of the PrO posterior. For further details, see Appendix D.6.
The results of our experiments are summarised in Figure 10. Here, we show the one-step-ahead predictive means over the period of November 2002 to March 2003 in the left column. We find that the richer parameter uncertainty conferred by the PrO posterior is able to handle the variability in the data more easily than the Gibbs posterior. This is also evident from the right column of the Figure, which charts the cumulative performance of the leave-one-out predictions. In summary, despite violating the assumption of independently sampled data, we can recover the clear benefits of PrO posteriors. This suggests that the independence assumptions imposed in our work are not required to justify using the PrO posterior in practice.
Figure 10: River flow data. In the left column we show the water flow (log scaled) in the winter leading up to the drought of 2003 along with the mean of the Gibbs and PrO predictive distributions respectively. In the right column we show the one-step-ahead cumulative mean predictive metrics.
5.3 Boston housing data
Strengthening the observations made in our previous experiment, we now demonstrate that PrO posteriors continue to function reliably in spatial settings by modelling the median house price across $n=506$ different census tracts in Boston in the 1970s and 1980s. These data were originally collected by Harrison and Rubinfeld (1978), and are summarised graphically in the left-most panel of Figure 11. Here, census tracts $i$ and $j$ are displayed as nodes in an undirected graph with adjacency matrix $\Omega$ whose entry at $(i,j)$ is given by $\omega_{i,j}=0$ unless the tracts $i$ and $j$ share at least one vertex, in which case $\omega_{i,j}=1$.
Median house values are then modelled as the conditional auto-regression
$y_{i}\mid y_{1},\dots y_{i-1},y_{i+1},\dots y_{n}\sim\mathsf{N}\left(\theta\sum_{j\neq i}\tilde{\omega}_{i,j}y_{j},\sigma^{2}\right),$
where $\tilde{\omega}_{i,j}$ denote the rescaled $\omega_{i,j}$ so that $\sum_{j\neq i}\tilde{\omega}_{i,j}=1$ for all $i$ and where $\sigma^{2}>0$ is a fixed constant which we estimate from data (for a more detailed explanation, see Appendix D.3). From this model, we construct two Gibbs and PrO posteriors for $\theta$ using the squared MMD and log score respectively; see Appendix D.8 for further details.
Importantly, the underlying data distribution is bimodal, and has its primary mode at around $\mathdollar 20,000$ and a second smaller mode at $\mathdollar 50,000$ (see Appendix D.8), with the second mode is due to censoring values exceeding $\mathdollar 50,000$ (Gilley and Pace, 1996). In other words, the data set has two sub-groups: tracts with median prices close to $\mathdollar 20,000$, and tracts with median prices are more than twice that value. Our results in Figure 11 show the effect on inference: while the Gibbs posterior fails to capture this hidden structure in the data, the PrO posterior transfers it to the parameter space. Its bi-modality reflects the fact that while there is a small group of census tracts with very high house prices surrounded by cheaper areas, a much larger group will be surrounded by similarly priced tracts. While the former group is accounted for through a smaller mode around $\theta\in[1.2,1.4]$, the latter is captured by the larger mode around $\theta\in[0.8,1.0]$. As the rightmost panel of Figure 11 confirms, this is not only a richer description of parameter uncertainty, but also leads to better predictive performance. Confirming our earlier observation that the choice of scoring rule seems to matter less for PrO posteriors, these results and interpretations are qualitatively the same whether one uses the squared MMD or the logarithmic scoring rule.
Figure 11: Top left: illustration of the Boston housing data set. Nodes correspond to census tracts, and shade represents price level. Census tracts are connected by edges if they are spatially adjacent. Middle columns: Gibbs and PrO parameter posteriors under the MMD score (top row) and the log score (bottom row). Right column: mean and standard error of the leave-one-node-out negative log likelihood (lower is better).
6 Computation via Wasserstein Gradient Flows
The tools for the computation of PrO posteriors are limited. In fact, most previous work targeting related objects did not even consider the possibility of exact computation, and instead focused on approximating $Q_{n}$ with a parameterised variational family (see Jankowiak et al., 2020a, b; Masegosa, 2020; Morningstar et al., 2022; Lai and Yao, 2024). The only pre-existing work directly targeting exact computation is Shen et al. (2025), which built on the ideas in Wild et al. (2023) and the mean-field Langevin algorithm of Del Moral (2013) to derive an asymptotically exact sampler for the special case where the scoring rule is the squared MMD. Here, we explain how this can be generalised to any PrO posterior.
We may rewrite the objective minimised in (4) as
                                                                                                    $\mathscr{L}(Q)$ ${}={}$ $\frac{\lambda_{n}}{n}\sum_{i=1}^{n}S\left(P_{Q},x_{i}\right)-\int\log\mathsf{d}\Pi(\theta)\,\mathsf{d}Q(\theta)+\int\log\mathsf{d}Q(\theta)\,\mathsf{d}Q(\theta).$                                                                                                      (14)
For functionals like this, the Wasserstein gradient flow (WGF) can be used to construct a gradient descent scheme via interacting particle systems derived from a McKean-Vlasov process. If the Wasserstein gradient of $\frac{1}{n}\sum_{i=1}^{n}S\left(P_{Q},x_{i}\right)$ is given by $\mathcal{W}(Q)$, the Wasserstein gradient of $\mathscr{L}(Q)$ with respect to $Q$ evaluated at $\vartheta\in\Theta$ is
                                                                                                    $\nabla_{\operatorname{W}}\mathscr{L}(Q)[\vartheta]$ ${}={}$ $\lambda_{n}\mathcal{W}(Q)[\vartheta]-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta)+\nabla_{\vartheta}\log\mathsf{d}Q(\vartheta).$                                                                                                      (15)
The WGF $\{Q_{t}\}_{t>0}$ defined by this gradient is the continuous-time process which evolves in the direction of steepest descent relative to the convex functional $Q\mapsto\mathscr{L}(Q)$ given by $\nabla_{\operatorname{W}}\mathscr{L}(Q)$. As $\mathscr{L}(Q)$ is convex, and as the minimiser of $\mathscr{L}(Q)$ is the PrO posterior, the WGF acts as a continuously-valued version of gradient descent on probability measures, so that $Q_{t}$ converges to the PrO posterior as $t\to\infty$ under mild conditions.
To implement this gradient descent algorithm numerically, we resort to the WGF’s mean-field representation, which gives rise to the stochastic differential equation (SDE)
                                                                                                    $\mathsf{d}\vartheta_{t}$ ${}={}$ $-\left\{\lambda_{n}\mathcal{W}(Q_{t})[\vartheta_{t}]-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta_{t})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t},\;\;\text{ for }\vartheta_{t}\sim Q_{t},$                                                                                                      (16)
and where the Brownian motion $B_{t}$ appears as the WGF of entropy corresponds to the heat equation (Jordan et al., 1998). Unfortunately, the above SDE still depends on the unknown measure $Q_{t}$, which makes exact forward simulation impossible. To remedy this, we use the approximations $Q_{t}^{(j)}(\theta)\approx\frac{1}{p}\sum_{\ell=1,\ell\neq j}^{p}\updelta_{\vartheta_{t}^{(\ell)}}$, where $\updelta_{x}$ denotes the Dirac measure at $x$, and $\vartheta_{t}^{(j)}$ for $j=1,2,\dots p$ are a set of $p$ interacting particles. This finally results in a system of $p$ interacting particles whose evolution is governed by the $p$ equations
                                                                                                    $\mathsf{d}\vartheta_{t}^{(j)}$ ${}={}$ $-\left\{\mathcal{W}(Q_{t}^{(j)})[\vartheta_{t}^{(j)}]-\nabla_{\vartheta}\log\mathsf{d}\Pi(\vartheta_{t}^{(j)})\right\}\mathsf{d}t+\sqrt{2}\mathsf{d}B_{t}^{(j)},\quad j=1,2\dots p,$                                                                                                      (17)
where $\{B_{t}^{(j)}:j=1,2,\dots p\}$ denotes $p$ independent Brownian motions. To obtain a high-fidelity approximation of the PrO posterior from this particle system, one discretises it into time steps $t_{1},t_{2},\dots$, and evolves it for a sufficiently long time. Afterwards, one averages over all time steps past an initial burn-in $\tau$ and approximates the PrO posterior as
                                                                                                    $Q_{n}\approx\frac{1}{|\{t_{i}:t_{i}>\tau\}|}\sum_{t_{i}>\tau}\widehat{Q}[t_{i}]\quad\text{ for }\quad\widehat{Q}[t_{i}]=\frac{1}{p}\sum_{j=1}^{p}\updelta_{\vartheta_{t_{i}}^{(j)}}.$                                                                                                      (18)
Here, the notation highlights that this approximation should be thought of as an average of $|\{t_{i}:t_{i}>\tau\}|$ measures $\widehat{Q}[t]$, each of which would recover $Q_{n}$ exactly if there were no discretisation error and both $p\to\infty$ and $t\to\infty$ (see e.g. Lemma 6 in Appendix E of Wild et al., 2023). We showcase this algorithm for the squared MMD and the logarithmic scoring rule in Appendix E.1, and elaborate on practical considerations in Appendix E.2.
@@ 7 Discussion
In this paper, we have advocated for a new statistical paradigm that defines epistemic uncertainty as a function of predictive capability. This led to PrO posteriors, which predictively dominate alternative methods for parameter inference like Gibbs posteriors and deliver interpretable and practically meaningful uncertainty quantification. We argued that PrO posteriors bridge the two main statistical traditions identified by Breiman (2001), and asserted that they lead to a version of epistemic uncertainty that is more expansive, adapts to model misspecification, and can avoid posterior collapse whenever this improves predictive performance. Throughout, we rigorously established these benefits and illustrated their practical importance on several instructive examples.
While we have presented evidence which advocates the adoption of PrO posteriors for inference and uncertainty quantification, much remains to be done. Perhaps the most pressing concern is the improvement of the available strategies for computation: sampling schemes obtained from interacting particle systems require further work before they can become attractive off-the-shelf algorithmic framework. Most notably, naïve implementations are infeasible for high-dimensional problems, generally computationally demanding, and require careful tuning of several hyperparameters. However, since research on improving such samplers has recently intensified (see Suzuki et al., 2023; Fu and Wilson, 2023; Liu et al., 2023; Chazal et al., 2025), we are hopeful that advances in the area will facilitate purpose-built algorithms which overcome these limitations.
Beyond computation, there are several outstanding theoretical questions that our current development has not explored yet. The most obvious of these concerns the choice of scoring rule. While our experiments and theory suggest that this choice is relatively unimportant in the case of convex recovery, we expect there to be important qualitative differences in behaviour for certain forms of non-trivial misspecification. For example, with Huber-contaminated data-generating processes of the form $P_{0}=(1-\varepsilon)P_{Q^{*}}+\varepsilon\cdot c$ for some $Q^{*}\in\mathcal{P}(\Theta)$ and some contamination $c\in\mathcal{P}(\mathcal{X})$, existing research on parametric inference (see e.g. Chérief-Abdellatif and Alquier, 2020, 2022; Dellaporta et al., 2022; Alquier et al., 2023; Alquier and Gerber, 2024) would indicate that PrO posteriors based on outlier-robust scoring rules like the squared maximum mean discrepancy should recover posteriors very similar to $Q^{*}$, while logarithmic scoring rules would behave completely differently. A second foundational question concerns the degree to which data can be allowed to be dependent: while our theory was derived under the assumption of independent data, our experiments seemed to suggest that the predictive benefits of PrO posteriors extend beyond this setting. Lastly, while it is clear that PrO posteriors are incoherent belief updates in the sense of Bissiri et al. (2016), it is worth investigating whether they can be re-cast as quasi-coherent updates on an extended space.
In the bigger picture, PrO posteriors have interesting connections to a number of other approaches grounded in frequentist statistics, and studying their intersection may result in mutually beneficial insights. For instance, the PrO posterior can be seen as a regularised frequentist estimate for the mixing measure in a nonparametric mixture model: most obviously, in the special case of the logarithmic scoring rule and when the prior regularisation term is dropped, (4) recovers nonparametric maximum likelihood exactly. Importantly, dropping the KL-term causes several practical issues that are absent from PrO posteriors, including computational difficulties due to non-identifiability (see e.g. Laird, 1978). Further, the corresponding measure-valued nonparametric maximum likelihood estimator generally cannot be interpreted as a tool for probabilistic uncertainty quantification—not only for conceptual reasons, but also because it is generally not unique and discrete; see Lindsay (1995, Theorem 21 in Chapter 5) and Jordan-Squire (2015).
Similarly, our work has clear parallels with the efforts of statistical learning theory to develop optimal aggregation procedures for predictors. For example, in his Saint-Flour lectures, Nemirovski (2000) introduces convex aggregation as the statistical task of finding the optimal convex combination of predictors. Intriguingly, for some model classes such as bounded regression with a finite number of candidates, the minimax-optimal convex aggregation procedure with respect to the squared risk can be shown to coincide precisely with (4) so long as the KL-regularisation term is removed (see, e.g., Lecué, 2013; Yang and Dunson, 2014). Though we have made no attempt at formalising these suspicions, this also suggests that PrO posteriors may be close to minimax-optimal convex aggregation procedures in certain problem classes.